OpenStudy (anonymous):

Find the exact value of the radical expression in simplest form. the square root of s to the eighth power + the square root of 25 times s to the eighth power + 2 the square root of s to the eighth power − the square root of s to the fourth power

OpenStudy (jdoe0001):

\(\bf \sqrt{s^8} + \sqrt{25s^8} + 2\sqrt{s^8} - \sqrt{s^4}\\\quad \\ \textit{keep in mind that}\\ s^8 = s^2 \times s^2 \times s^2 \times s^2\\ 5^2 = 25\)